FULL SCREtN PRINTER VERSION ㅖ BACK NEXT Rudd Clothiers is a small company that manufactures tall-mens suits. The company has used a standard cost accounting system. In May 2017, 10,100 suits were produced. The following standard and actual cost data applied to the month of May when normal capacity was 14,000 direct labor hours. All materials purchased were used. Standard (per unit) Actual Cost Element Direct materials Direct labor Overhead 9 yards at $4.00 per yard 1.10 hours at $14.00 per hour 1.10 hours at $6.80 per hour (fixed $4.00; variable $2.80) $352,275 for 91,500 yards ($3.85 per yard) $171,245 for 11,810 hours ($14.50 per hour) $49,800 fixed overhead $37,500 variable overhead Overhead is applied on the basis of direct labor hours. At normal capacity, budgeted fixed overhead costs were $56,000, anil budgeted variable overhead was $39,200. Compute the total, price, and quantity variances for (1) materials and (2) labor. (Round answers to O decimal places, e.g. 125.) (1) Total materials variance 13.725 Favorable Materials price variance 2400 Favorable Materials quantity variance 11325 Favorable (2) Total labor variance 5905 Favorable Labor price variance 9800 Favorable Labor quantity variance 5705 Favorable
Answer #1

Ans-(a) (1)-

Actual Production 10,100
Standard Quantity of Direct Material Required per unit 9
Standard Quantity of Direct Material Required (SQ) (10,100*9) 90,900
Standard Price per Yard (SP) 4
Actual Direct Material (AQ) 91,500
Actual Price per Pound (AP) 3.85

Total Material Variance = (SP*SQ)-(AP *AQ)

=(4*90,900)- (3.85*91,500)

=$363,600-$352,275

=$11,325 Favorable

Direct Material Price Variance = AQ* (SP-AP)

=91,500* (4-3.85)

=91,500*0.15

=$13,725 Favorable

Direct Material Quantity Variance= SP * (SQ-AQ)

=4* (90,900-91,500)

=4*-600

=-2,400 Unfavorable

(a) (2)-

Actual Production 10,100
Standard Hours Required per Unit 1.10
Standard Hours Requires (SH) (10,100*1.10) 11,110
Standard Rate per Hour (SR) 14
Actual Hours Required (AH) 11,810
Actual Rate per Hour (AR) 14.50

Total Labor Variance= (SR*SH)- (AR*AH)

=(14*11,110)- (14.50*11,810)

=$155,540- $171,245

=-15,705 Unfavorable

Direct Labor Rate Variance =AH * (SR-AR)

=11,810* (14-14.50)

=11,810*-0.50

=-$5,905 Unfavorable

Direct Labor Efficiency Variance= SR* (SH-AH)

=14*( 11,110-11,810)

=14*-700

=-9,800 Unfavorable

Kindly give thumbsup if u like my answer...Thanks!!!